OpenStudy (anonymous):

Which ordered pair is on the graph of the equation 2x + 5y = 4? A. (3, –2) B. (0, 1) C. (0, –1) D. (7, –2)

OpenStudy (nincompoop):

plug in the numbers and see

OpenStudy (anonymous):

I DONT GET GRAPHS I'VE BEEN WORKING FOR 6 HOURS STRAIGHT COULD YOU PLEASE JUST HELP ME?!

OpenStudy (nincompoop):

there is no graph in that problem all you have to do is plug the values into x and y and then see if the equation matches

OpenStudy (anonymous):

YOUR RIGHT I SAID THAT TOO!!!!

OpenStudy (nincompoop):

so start plugging the numbers into the x and y

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=2x+%2B+5y+%3D+4

OpenStudy (anonymous):

The link but it still doesn't make sense.

OpenStudy (nincompoop):

(3, –2) 2(3) + 5(-2) = 4 TRUE OR FALSE

OpenStudy (nincompoop):

that's all you gotta do, test all the ordered pairs and see

OpenStudy (nincompoop):

you don't even need wolframalpha for this kind of problem

OpenStudy (anonymous):

I don't know I think the rest are right so..

OpenStudy (nincompoop):

do you know how to multiply, add, subtract and divide?

OpenStudy (anonymous):

Yes but i am tired of working all day....

OpenStudy (nincompoop):

(3, –2) 2(3) + 5(-2) = 4 6 - 10 = 4 -4 = 4 TRUE OR FALSE?

OpenStudy (nincompoop):

well then go get your rest and continue tomorrow

OpenStudy (anonymous):

I have to do this tonight right NOW due

OpenStudy (nincompoop):

then do it I showed you how to do it